# Fast and Accurate Charge Transfer Excitations via Nested Aufbau Suppressed Coupled Cluster

### v1

@@ Abstract
Modeling charge transfer well can require treating post-excitation orbital relaxations and handling medium to large molecules in realistic environments. By combining a state-specific correlation treatment with such orbital relaxations, Aufbau suppressed coupled cluster has proven accurate for charge transfer, but, like many coupled cluster methods, it struggles with large system sizes. We derive a low-cost Aufbau suppressed second order perturbation theory and show that, by nesting a small coupled cluster treatment inside of it, computational cost and scaling are reduced while accuracy is maintained. Formal asymptotic costs are dropped from iterative $N^{6}$ to non-iterative $N^{5}$ plus iterative $N^{3}$, and we test an initial implementation that can handle about 100 atoms and 800 orbitals on a single computational node. Charge transfer excitation energy errors are below 0.1 eV on average, a 0.25 eV improvement over $N^{6}$-cost equation of motion coupled cluster with singles and doubles.
@@
ASCC and its partially linearized cousin (PLASCC) have a two-fold advantage in handling post-excitation orbital relaxations that makes them particularly well positioned to model CT. [20, 21] First, they can be constructed on top of a reference wave function that already captures orbital relaxation effects at the mean field level. Second, their full $\hat{T}_{1}$ singles operators allow them to fine tune orbital relaxations in the presence of correlation effects.[22] The ability to deliver a CCSD-like weak correlation treatment alongside and informed by these orbital relaxations has allowed PLASCC to deliver $\sim$0.1 eV excitation energy errors across a wide variety of excited states, including Rydberg and CT states. [21] While this accuracy is especially noteworthy for CT, where EOM-CCSD’s errors are $\sim$0.35 eV, the $N^{6}$ cost scaling with system size $N$ has so far limited the applicability of ASCC and PLASCC to relatively small systems.
@@
After perturbatively partitioning the Hamiltonian (see Supporting Information), its double similarity transform can also be partitioned by perturbative order.
@@
While the working equations for this PT are a bit more complicated than for MP2 due to the presence of $\hat{T}^{(0)}$, they do closely mirror MP2’s computational cost. Once the two-electron integrals are available in the molecular orbital (MO) basis, every term in the working equations scales as $O(N^{4})$ or less, as is the case for MP2. Even better, the amplitude equations in Eqn. 11 are linear in the highest order terms — which at each order are the ones being solved for — and are small-block block-diagonal within a semicanonical basis, allowing them to be solved in a non-iterative manner (see Supporting Information). Therefore, the headline scaling of this PT, in exact asymptotic equivalence to MP2, comes from the non-iterative $O(N^{5})$ cost of transforming the two-electron integrals into the MO basis. Compared to the iterative $O(N^{6})$ scaling of ASCCSD, this is a substantial improvement in computational efficiency and notably avoids some of the most memory intensive terms found in CCSD and ASCCSD. Even when compared to relatively affordable wave function correlation methods for excited states, such as the iterative $O(N^{5})$ scaling CC2[28] and ADC(2)[29, 30, 31] methods, the non-iterative $O(N^{5})$ scaling of a routine integral transformation represents a significant cost reduction.
@@
Figure 1: Excitation energy error distributions for various methods on 130 valence and Rydberg single-CSF singlet states from the QUEST benchmark in an aug-cc-pVDZ basis. For the PT alone, 17 outlier states with a $<-1.0$ eV error are not shown. Note the change of scale in the y-axis between the top and bottom rows. The mean signed error (MSE), mean unsigned error (MUE), and standard deviations (SD) are shown in insets. Reference values are of at least of EOM-CCSDT quality.
@@
Figure 2: Mean unsigned errors (MUEs) for various methods on 7 CT excitations coming mostly from the benchmark of Szalay and coworkers in a cc-pVDZ basis. MUEs extending beyond the plot are printed at the top of each bar. Reference values are EOM-CCSDT except for one LR-CC3 value.
@@
Before we go about nesting a small ASCC treatment within this PT, it is worthwhile to evaluate its accuracy as a standalone method. To do so, we test the PT on a collection of 130 singly excited, single CSF, singlet valence and Rydberg states from the QUEST molecular benchmark[32, 33] as well as a set of 7 CT states mainly originating from the benchmark of Szalay and coworkers.[14] The results are summarized in Figures 1 and 2, with additional details available in the Supporting Information. While computationally efficient, the PT is unsurprisingly less accurate than its parent CC theory. Though it performs well in many cases, it produces a substantial number of severe errors, some larger than an eV, particularly in the aromatic molecules where PLASCC makes the largest improvements over ASCC. Indeed, this pattern helps corroborate the original motivation for PLASCC, which was that in these systems the perturbation is large enough that higher order and poorly balanced terms in ASCC become important. [21] At the very least, these PT results suggest that the perturbation in ASCC theory is indeed larger in many of these ring states.
@@
Of particular concern for the present study is that the PT fares worse in CT states (errors of $\sim$0.6 eV) than it does in valence and Rydberg states (errors of $\sim$0.4 eV). This worsening contrasts sharply with PLASCC, which produces errors of $\sim$0.1 eV in each category. One culprit is likely the truncation of singles contributions to low orders in PT theory, which interferes with the parent theory’s ability to provide CC-singles-style corrections to the initial mean field orbital relaxations. [20, 21] Another factor is that more electron pairs experience large, excitation-induced changes to their electron correlation in CT states, because CT excitations make major changes to how close various electrons are to each other. It is to be expected that second order PT would be less effective at capturing these changes than a CCSD-style approach. These challenges notwithstanding, the PT does offer significantly higher accuracy for CT than hybrid TD-DFT methods, even when long-range-corrected functionals like $\omega$B97X-D [34] and CAM-B3LYP [35] are used.
@@
We therefore begin by performing full MP2 and Aufbau suppressed PT calculations on the ground and excited states, respectively. While these use (semi)canonical bases to ensure a non-iterative $O(N^{5})$ cost, we then proceed to localize and match orbitals between the ground and excited state (see Supporting Information) in preparation for the correlation analysis. In this matched local basis, we construct ground and excited state measures for the correlation contributions related to each orbital. Specifically, for a given orbital, we add together all of the PT’s amplitude contributions to Eq. 9 in which the amplitude involved has at least one index on that orbital. We then compare this correlation measure between the ground and the excited state, and, if they differ by more than a small set threshold, that orbital gets flagged as needing to be part of the nested CC treatment. One can imagine other screening approaches, and it’s hard to know what the optimal approach would be, but we will emphasize some strong formal properties of this approach: it preserves size consistency, extensivity, and intensivity and, for a given localized orbital, is invariant to rotations among the other localized occupied or virtual orbitals.
@@
Turning now to the CT results in Figure 2, we see a similarly encouraging story unfold. Nested ASCC and PLASCC are much more accurate than the PT theory on its own. PLASCC and nested PLASCC are again more accurate than ASCC, although in this case nested ASCC is not an improvement over full ASCC. Comparing to other excited state theories, we see that nested PLASCC maintains PLASCC’s 0.25 eV improvement over EOM-CCSD, which is now especially significant in light of the non-iterative $O(N^{5})$ vs iterative $O(N^{6})$ computational costs. Nested PLASCC far outstrips the TD-DFT accuracy of the density functionals we test here, and is also a notable improvement over the much wider range of TD-DFT methods tested by Mester and Kállay. [13] Seeing that PLASCC’s accuracy is maintained under nesting in valence, Rydberg, and even CT states, we now turn our attention to evaluating computational efficiency in larger systems where tidy theoretical energy benchmarks are not available.
@@
We begin with the simple gas phase valence excitation in Figure 3A to demonstrate a few key points. First, this excitation is an example where the nested CC refinement makes a significant, $\sim$0.15 eV difference compared to the PT alone. This difference is likely driven by the importance of post-excitation orbital relaxations on the relatively polarizable sulfur atom following the delocalization of the lone pair electron into the $\pi^{*}$ system. Second, despite this delocalization, the excitation remains localized compared to the extent of the molecule, which should leave most orbitals’ correlation contributions mostly unaffected. Finally, its computational cost grows with system size as we would expect it to, as seen in Figure 4.
@@
For an explicit solvation example, we examine in Figure 3B the $n$ to $\pi^{*}$ transition in solvated acetone. The explicit waters play a non-innocent role, as the hydrogen bonding that initially stabilizes the ground state is disrupted when one of the hydrogen-bond-accepting lone pair electrons is transferred into the $\pi^{*}$ orbital by the excitation. Crucially, despite the strong solute-solvent interactions and the three-dimensional nature of this example, the orbitals flagged for CC refinement remain few in number and are tightly localized in space. We find it especially noteworthy that, despite no explicit input to this effect from the user, the PT’s correlation analysis selected CC orbitals in a physically intuitive manner: the water orbitals it picks out are the ones that had their strong hydrogen bonding interactions perturbed by the excitation.
Finally, we examine the donor-bridge-acceptor CT example of Figure 3C, which results from adding an alkane bridge to one of our earlier CT tests. In contrast to the last two examples in which the excitation was localized in only one region of space, here the PT’s correlation analysis flags two separate regions for CC treatment: one on the donor and one on the acceptor, again following what physical intuition would suggest. While this donor/acceptor effect leads to a larger number of orbitals in the CC calculation, this number stops changing with system size once the bridge reaches just three carbons in length. Thus, regardless of whether the excitation itself is a local valence excitation or a non-local CT, we see that the nesting approach successfully demarcates a physically intuitive set of orbitals for careful CC treatment while dramatically reducing the computational cost compared to a full-system CC evaluation.
@@ I Computational Methods
Our EOM-CCSD calculations and Foster-Boys orbital localizations[36] were performed with PySCF,[37, 38, 39] our DFT and EOM-CCSDT calculations were performed with Q-Chem,[40] and our LR-CC3 calculations were performed with PSI4.[41] Though the literature and reference calculations were performed utilizing the frozen core approximation, our EOM-CCSD, PT, and ASCC calculations were performed without this approximation. Nevertheless, this is expected to make only a small difference to the excitation energies ($\sim$0.02 eV)[32, 33] and thus the high accuracy frozen-core values still make excellent references. The geometries for the large molecule valence and charge transfer excitations were determined via MP2/cc-pVDZ geometry optimizations in Q-Chem, while the solvated acetone geometry was determined using the solvator method in ORCA.[42, 43] Explicit geometries can be found in the Supporting Information.
For the excited state PT, ASCC, and PLASCC calculations, the excited state reference was determined via the excited state mean field (ESMF) method,[44, 45, 46] which can be summarized as a CIS wavefunction with full orbital relaxations. Any CSF in the ESMF reference with a singular value greater than 0.2 was considered part of the excited state reference. Block-sparse [47] Cholesky decompositions [48] with a threshold of $10^{-7}$ Hartree were used to evaluate MO basis two electron integrals, with Libint[49] supplying the underlying atomic orbital integrals. For the ASCC based approaches, convergence was considered achieved when the maximum residual was no greater than $10^{-5}$ and the energy changed by less than $10^{-7}$ Hartree in an iteration. For all other methods, default convergence criteria were used. On rare occasions, convergence difficulties were encountered for ASCC based approaches. Details regarding these states may be found in the Supporting Information. For the nested ASCC calculations, three orbital energy thresholds for establishing the small orbital set for the CC part were tested. The accuracy data in the main text are for a 0.005 eV threshold, with additional accuracy data for 0.01 eV and 0.0025 eV thresholds available in the SI. The scaling analysis was performed using the 0.01 eV threshold.

### v2

@@ Abstract
Modeling charge transfer well can require treating post-excitation orbital relaxations and handling medium to large molecules in realistic environments. By combining a state-specific correlation treatment with such orbital relaxations, Aufbau suppressed coupled cluster has proven accurate for charge transfer, but, like many coupled cluster methods, it struggles with large system sizes. We derive a low-cost Aufbau suppressed second order perturbation theory and show that, by nesting a small coupled cluster treatment inside of it, computational cost and scaling are reduced while accuracy is maintained. Formal asymptotic costs are dropped from iterative $N^{6}$ to non-iterative $N^{5}$ plus iterative $N^{3}$, and we test an initial implementation that can handle about 100 atoms and 800 orbitals on a single computational node. Charge transfer excitation energy errors are typically below 0.1 eV, with an average 0.25 eV improvement over $N^{6}$-cost equation of motion coupled cluster with singles and doubles.
@@
ASCC and its partially linearized cousin (PLASCC) have a two-fold advantage in handling post-excitation orbital relaxations that makes them particularly well positioned to model CT. [20, 21] First, they can be constructed on top of a reference wave function that already captures orbital relaxation effects at the mean field level. Second, their full $\hat{T}_{1}$ singles operators allow them to fine tune orbital relaxations in the presence of correlation effects.[22] The ability to deliver a CCSD-like weak correlation treatment alongside and informed by these orbital relaxations has allowed PLASCC to deliver $\sim$0.1 eV excitation energy errors across a wide variety of excited states, including Rydberg and CT states. [21] While this accuracy is especially noteworthy for CT, where EOM-CCSD’s errors are $\sim$0.375 eV, the $N^{6}$ cost scaling with system size $N$ has so far limited the applicability of ASCC and PLASCC to relatively small systems.
@@
After perturbatively partitioning the Hamiltonian — see the Supporting Information (SI) — its double similarity transform can also be partitioned by perturbative order.
@@
While the working equations for this PT are a bit more complicated than for MP2 due to the presence of $\hat{T}^{(0)}$, they do closely mirror MP2’s computational cost. Once the two-electron integrals are available in the molecular orbital (MO) basis, every term in the working equations scales as $O(N^{4})$ or less, as is the case for MP2. Even better, the amplitude equations in Eqn. 11 are linear in the highest order terms — which at each order are the ones being solved for — and are small-block block-diagonal within a semicanonical basis, allowing them to be solved in a non-iterative manner (see SI). Therefore, the headline scaling of this PT, in exact asymptotic equivalence to MP2, comes from the non-iterative $O(N^{5})$ cost of transforming the two-electron integrals into the MO basis. Compared to the iterative $O(N^{6})$ scaling of ASCCSD, this is a substantial improvement in computational efficiency and notably avoids some of the most memory intensive terms found in CCSD and ASCCSD. Even when compared to relatively affordable wave function correlation methods for excited states, such as the iterative $O(N^{5})$ scaling CC2[28] and ADC(2)[29, 30, 31] methods, the non-iterative $O(N^{5})$ scaling of a routine integral transformation represents a significant cost reduction.
@@
Figure 1: Excitation energy error distributions for various methods on 130 valence and Rydberg single-CSF singlet states from the QUEST benchmark in an aug-cc-pVDZ basis. For the PT alone, 17 outlier states with a $<-1.0$ eV error are not shown. Note the change of scale in the y-axis between the top and bottom rows. The mean signed error (MSE), mean unsigned error (MUE), and standard deviations (SD) are shown in insets. Reference values are of at least EOM-CCSDT quality.
@@
Figure 2: Mean unsigned errors (MUEs) for various methods on 16 intermolecular and intramolecular CT excitations in a cc-pVDZ basis. MUEs extending beyond the plot are printed at the top of each bar. Reference values are EOM-CCSDT except for one LR-CC3 value.
@@
Before we go about nesting a small ASCC treatment within this PT, it is worthwhile to evaluate its accuracy as a standalone method. To do so, we test the PT on a collection of 130 singly excited, single CSF, singlet valence and Rydberg states from the QUEST molecular benchmark[32, 33] as well as a set of 16 CT states, with intermolecular CT states coming from the benchmark of Szalay and coworkers[14] and intramolecular CT states coming mainly from the long range CT states in the benchmark of Loos and coworkers.[34] The results are summarized in Figures 1 and 2, with additional details available in the SI (note that we will develop the nested theories referred to in these figures further down in the main text). While computationally efficient, the PT is unsurprisingly less accurate than its parent CC theory. Though it performs well in many cases, it produces a substantial number of severe errors, some larger than an eV, particularly in the aromatic molecules where PLASCC makes the largest improvements over ASCC. Indeed, this pattern helps corroborate the original motivation for PLASCC, which was that in these systems the perturbation is large enough that higher order and poorly balanced terms in ASCC become important. [21] At the very least, these PT results suggest that the perturbation in ASCC theory is indeed larger in many of these ring states.
@@
Of particular concern for the present study is that the PT fares worse in CT states (errors of $\sim$0.5 eV) than it does in valence and Rydberg states (errors of $\sim$0.4 eV). This worsening contrasts sharply with PLASCC, which typically produces errors of $\sim$0.1 eV in each category. One culprit is likely the truncation of singles contributions to low orders in PT theory, which interferes with the parent theory’s ability to provide CC-singles-style corrections to the initial mean field orbital relaxations. [20, 21] Another factor is that more electron pairs experience large, excitation-induced changes to their electron correlation in CT states, because CT excitations make major changes to how close various electrons are to each other. It is to be expected that second order PT would be less effective at capturing these changes than a CCSD-style approach. These challenges notwithstanding, the PT does offer significantly higher accuracy for CT than hybrid TD-DFT methods, even when long-range-corrected functionals like $\omega$B97X-D [35] and CAM-B3LYP [36] are used.
@@
We therefore begin by performing full MP2 and Aufbau suppressed PT calculations on the ground and excited states, respectively. While these use (semi)canonical bases to ensure a non-iterative $O(N^{5})$ cost, we then proceed to localize and match orbitals between the ground and excited state (see SI) in preparation for the correlation analysis. (While for simplicity we employ standard Foster-Boys orbital localization [37] in this initial investigation, related approaches [38, 39, 40, 41, 42, 43, 44, 45, 46, 47] suggest that localizations of subsets of the MP2 or Aufbau suppressed PT natural orbitals [48] or natural transition orbitals [49] could be even more effective, especially for the virtual orbitals.) In this matched local basis, we construct ground and excited state measures for the correlation contributions related to each orbital. Specifically, for a given orbital, we add together all of the PT’s amplitude contributions to Eq. 9 in which the amplitude involved has at least one index on that orbital. We then compare this correlation measure between the ground and the excited state, and, if they differ by more than a small set threshold, that orbital gets flagged as needing to be part of the nested CC treatment. One can imagine other screening approaches, and it’s hard to know what the optimal approach would be, but we will emphasize some strong formal properties of this approach: it preserves size consistency, extensivity, and intensivity and, for a given localized orbital, is invariant to rotations among the other localized occupied or virtual orbitals.
@@
Figure 3: Tests on the ${}^{1}A^{\prime}$ (dashed lines) and ${}^{1}A^{\prime\prime}$ (solid lines) $n\rightarrow\pi^{*}$ CT states as a water is moved near its hydrogen bond acceptor. Shown at the top (a) is a sketch of the water’s positions at geometry numbers 0, 2, and 4, from left to right, respectively. The three plots show (b) excitation energy errors relative to CC3, (c) the excited state potential energy surfaces, whose zeros of energy are set to geometry 2 and where the $A^{\prime\prime}$ surfaces have then been shifted up by 2 kcal/mol for clarity, and (d) the weight of the largest non-CT component in the state. The basis is cc-pVDZ. See text and SI for details.
While our nesting approach and double similarity transform contain parallels to multilevel CC (MLCC) theory [50] and similarity transformed EOM (STEOM), [51] respectively, it is important to appreciate the differences. MLCC, like EOM-CCSD, employs the equation-of-motion formalism for excited states, [50, 45, 46] and so it will inherit EOM-CCSD’s difficulties with post-excitation orbital relaxations. ASCC and PLASCC, in contrast, provide a more robust orbital relaxation treatment, as discussed above. This difference explains their accuracy advantage over EOM-CCSD in CT states, an advantage that we will see is maintained even when nested inside the new PT. In STEOM, electron affinity and ionization potential calculations are used to construct the second similarity transform, [51] which for long-range CT should build in the appropriate orbital relaxations. However, it is less obvious that the same will be true for local excitations, where no region of the molecule loses or gains an electron. In ASCC, the second similarity transform is constructed from an orbital-relaxed, neutral excited state reference, and any further fine tuning of the relaxation can be done state-specifically via the exponentiated singles operator.
@@
Turning now to the CT results in Figure 2, we see a similarly encouraging story unfold. Nested ASCC and PLASCC are much more accurate than the PT theory on its own. PLASCC and nested PLASCC are again more accurate than ASCC, although in this case nested ASCC is not a significant improvement over full ASCC. For all but two states, nested PLASCC closely maintains PLASCC’s accuracy, ultimately resulting in a 0.25 eV improvement over EOM-CCSD, which is now especially significant in light of the non-iterative $O(N^{5})$ vs iterative $O(N^{6})$ computational costs. Nested PLASCC far outstrips the TD-DFT accuracy of the density functionals we test here, and is also a notable improvement over the much wider range of TD-DFT methods tested by Mester and Kállay. [13] With these initial surveys confirming nested PLASCC’s accuracy in valence, Rydberg, and CT states, we will now explore an example at the intersection of CT and hydrogen bonding.
In the system shown in Figure 3(a), a water molecule is moved past its low-energy hydrogen bonding arrangement with the acceptor moiety’s oxygen atom. This displacement test, the geometries for which are taken from another recent PT study, [52] allows us to examine the behavior of different methods’ excitation energies, potential energy surfaces (PESs), and state characters as the local environment of a CT acceptor is modified by a non-innocent solvent effect. We look in particular at the ${}^{1}A^{\prime}$ and ${}^{1}A^{\prime\prime}$ CT states that promote from the nitrogen lone pair to the in-plane and out-of-plane $\pi^{*}$ orbitals of the CCO group, respectively. The remaining panels of Figure 3 reveal that, among the three methods tested against CC3, nested PLASCC is the only method that simultaneously offers reasonably accurate excitation energies, potential energy surfaces within 1 kcal/mol of CC3, and accurate state characters for both states at all five geometries. As usual, TD-DFT severely underestimates the CT excitation energies, although it does prove quite accurate for the excited state PESs. The latter observation is somewhat surprising in light of the fact that it erroneously mixes a large portion of $\pi\rightarrow$Rydberg character into the $A^{\prime\prime}$ state at some geometries, seemingly a result of its too-low excitation energy causing two states to appear nearly degenerate when in fact they are not. Similarly, EOM-CCSD incorrectly mixes a large Rydberg component into the $A^{\prime}$ state at all geometries. In that case, however, the erroneous state character is sufficient to significantly alter the PES, which underestimates the energy needed to move the water from geometry 2 to geometry 4 by more than 1 kcal/mol. This underestimation makes sense, as, with half the state not putting negative charge on the acceptor, it becomes too easy to pull the partially positive hydrogen atom away from it. Although this is only one test, it does suggest that PES explorations may be a useful application of nested PLASCC once its analytic gradients have been implemented. For now, we will turn our attention to a more pressing concern, which is whether the theoretically expected cost scaling of the approach is seen in practice.
@@
We begin with the simple gas phase valence $n\rightarrow\pi^{*}$ excitation in a thiopropynal-based molecule with an extended alkane chain (Figure 4A) to demonstrate a few key points. First, this excitation is an example where the nested CC refinement makes a significant, $\sim$0.15 eV difference compared to the PT alone. This difference is likely driven by the importance of post-excitation orbital relaxations on the relatively polarizable sulfur atom following the delocalization of the lone pair electron into the $\pi^{*}$ system. Second, despite this delocalization, the excitation remains localized compared to the extent of the molecule, which should leave most orbitals’ correlation contributions mostly unaffected. Finally, its computational cost grows with system size as we would expect it to, as seen in Figure 5.
@@
For an explicit solvation example, we examine in Figure 4B the $n\rightarrow\pi^{*}$ transition in solvated acetone. The explicit waters play a non-innocent role, as the hydrogen bonding that initially stabilizes the ground state is disrupted when one of the hydrogen-bond-accepting lone pair electrons is transferred into the $\pi^{*}$ orbital by the excitation. Crucially, despite the strong solute-solvent interactions and the three-dimensional nature of this example, the orbitals flagged for CC refinement remain few in number and are tightly localized in space. We find it especially noteworthy that, despite no explicit input to this effect from the user, the PT’s correlation analysis selected CC orbitals in a physically intuitive manner: the water orbitals it picks out are the ones that had their strong hydrogen bonding interactions perturbed by the excitation.
Finally, we examine the donor-bridge-acceptor CT example of Figure 4C, which results from adding an alkane bridge to one of our earlier CT tests. In this state, a lone pair electron on the nitrogen donor is transferred to the $\pi^{*}$ system on the opposite end of the molecule. In contrast to the last two examples in which the excitation was localized in only one region of space, here the PT’s correlation analysis flags two separate regions for CC treatment: one on the donor and one on the acceptor, again following what physical intuition would suggest. While this donor/acceptor effect leads to a larger number of orbitals in the CC calculation, this number stops changing with system size once the bridge reaches just three carbons in length. Thus, regardless of whether the excitation itself is a local valence excitation or a non-local CT, we see that the nesting approach successfully demarcates a physically intuitive set of orbitals for careful CC treatment while dramatically reducing the computational cost compared to a full-system CC evaluation.
@@ I Computational Methods
Our EOM-CCSD calculations and Foster-Boys orbital localizations[37] were performed with PySCF,[53, 54, 55] our DFT and EOM-CCSDT calculations were performed with Q-Chem,[56] and our LR-CC3 calculations were performed with PSI4.[57] Though the literature and reference calculations were performed utilizing the frozen core approximation, our EOM-CCSD, PT, and ASCC calculations were performed without this approximation. Nevertheless, this is expected to make only a small difference to the excitation energies ($\sim$0.02 eV)[32, 33] and thus the high accuracy frozen-core values still make excellent references. The geometries for the large molecule valence and charge transfer excitations were determined via MP2/cc-pVDZ geometry optimizations in Q-Chem, while the solvated acetone geometry was determined using the solvator method in ORCA.[58, 59] Explicit geometries can be found in the SI.
For the excited state PT, ASCC, and PLASCC calculations, the excited state reference was determined via the excited state mean field (ESMF) method,[60, 61, 62] which can be summarized as a CIS wavefunction with full orbital relaxations. Any CSF in the ESMF reference with a singular value greater than 0.2 was considered part of the excited state reference. Block-sparse [63] Cholesky decompositions [64] with a threshold of $10^{-7}$ Hartree were used to evaluate MO basis two electron integrals, with Libint[65] supplying the underlying atomic orbital integrals. For the ASCC based approaches, convergence was considered achieved when the maximum residual was no greater than $10^{-5}$ and the energy changed by less than $10^{-7}$ Hartree in an iteration. For all other methods, default convergence criteria were used. On rare occasions, convergence difficulties were encountered for ASCC based approaches. Details regarding these states may be found in the SI.
For the nested ASCC calculations, three orbital energy thresholds for establishing the small orbital set for the CC part were tested. The benchmark data in the main text are with a 0.005 eV threshold, with additional accuracy data for 0.01 eV and 0.0025 eV thresholds available in the SI. For the test in Figure 3, the number of CC occupied and virtual orbitals was fixed with the tighter 0.0025 eV threshold at the equilibrium geometry and enforced across all other geometries on the surface in order to promote smoothness, and the excited state absolute energy was determined by adding the PT2 absolute energy of the ground state to the nested PLASCC excitation energy. Finally, the scaling analysis was performed using the 0.01 eV threshold.